OpenStudy (anonymous):

.. i need help finding the geometric mean of [4\sqrt{3}\] and [10\sqrt{3}\]

hero (hero):

Hint: To find the geometric mean of two numbers, a and b, set up the following proportion: \(\large\frac{a}{x} = \frac{x}{b}\)

hero (hero):

Then cross multiply. In this case a = \(4\sqrt{3}\) b = \(10\sqrt{3}\)

OpenStudy (anonymous):

but i still dont understand it.

hero (hero):

You don't know how to substitute numerical values for variables into the given formula?

OpenStudy (anonymous):

yes i have that part

hero (hero):

Okay, so you don't know how to cross multiply?

OpenStudy (anonymous):

i got the answer . thank you

hero (hero):

You got the answer? How?

hero (hero):

Did you do it yourself or did you use some other program? I suspect you used some other program. You're too confident in your answer, which you didn't even consider asking me whether or not it was correct.

OpenStudy (anonymous):

i did it my self i got 10.95 rounded up to 11.0

OpenStudy (anonymous):

it came to x^2=120 the square root of 120 is 10.954

hero (hero):

They gave you the problem in radicals, so your result should also be in radicals.

OpenStudy (anonymous):

2sqrt 30